How many grams of lead nitrate are needed to produce 105.25 g of sodium nitrate?
1
Expert's answer
2020-03-09T09:45:14-0400

"Pb(NO_3)_2 + 2NaOH = 2NaNO_3 + Pb(OH) _2"

"\nu(NaNO_3) = m/M = 105.25/85 = 1.24 mol."

1.24 : 2 = "\nu(Pb(NO_3)_2)" : 1

"\nu(Pb(NO_3)_2)=1.24/2=0.62 mol." "m(Pb(NO_3)_2)=\nu*M=0.62*331=205.22 g."


Answer: 205.22 g.